Newtown Propane currently has only a wholesale division and uses only equity capital; however, it is considering creating marketing and retail divisions. Its beta is currently 1.3. The marketing division is expected to have a beta of 2.1, because it will have more risk than the firm’s wholesale division. The retail division is expected to have a beta of 0.6, because it will have less risk than the firm’s wholesale division. The risk-free rate is 4.4%, and the market risk premium is 6.7%. Based on this information, fill in the missing cost of capital information below:

Wholesale division     

13.11%

8.80%

10.12%

3.08%

Marketing division     

18.47%

20.97%

19.82%

19.42%

Retail division     

20.42%

20.52%

19.22%

8.42%

If 55% of Newtown Propane’s total value ends up in the wholesale division, 20% in the marketing division, and 25% in the retail division, then its investors should require a return of   .

17.79%

14.34%

13.04%

15.89%
Answer #1

The cost of capital for the Wholesale division as per the CAPM model is:

Re = Rf + beta ( Rm - Rf)

= 4.4 + 1.3* 6.7

=,13.11%

So, the correct option is option 1.

The cost of capital for Marketing division is :

= 4.4 + 2.1* 6.7

=18.47%

So, the correct option is option 1.

The cost of capital for the retail division is :

= 4.4 + 0.6* 6.7

= 8.42%

So, the correct option is option D.

The cost of capital will be :

0.55*0.1311 + 0.2*0.1847 + 0.25* 0.0842

= 0.072105 + 0.03694 + 0.02105

= 13.01%

So, the correct option is option C.